OpenStudy (anonymous):

a company offers a starting yearly salary of $33,000 with raises of 2,000 per year.find the total salary over a Ten-year period. the TOTal salary is__.

OpenStudy (anonymous):

pl help me

OpenStudy (anonymous):

im lost but i tink it is 350,000

OpenStudy (gorv):

raise after 10 year= 2000*10=20000 salary=20000+33000=53000

OpenStudy (gorv):

it will form a aritmatic progression

OpenStudy (anonymous):

53000

OpenStudy (gorv):

33000,35000,................53000 a=33000 common difference=2000

OpenStudy (anonymous):

or 53,000

OpenStudy (gorv):

@sexybadass22 53000 is his salary of one year not total salary

OpenStudy (gorv):

are asking about sum of salary of all the years???

OpenStudy (anonymous):

53000-10

OpenStudy (anonymous):

530,000 for 10 year

OpenStudy (anonymous):

im right

OpenStudy (anonymous):

1st year is only $33,000 2nd year is 33,000 + 2000 3rd year is 33,000 + 4000 4th year is 33,000 + 6000 5th year is 33,000 + 8000 6th year is 33,000 + 10,000 7th year is 33,000 + 12,000 8th year is 33,000 + 14,000 9th year is 33,000 + 16,000 10th year is 33,000 + 18,000 So that makes 10* 33,000 + 200,000 or $530,000